OpenStudy (anonymous):

Solve this system of equations using elimination. -3x-4y=7 -2x+4y=3 x= y=

OpenStudy (anonymous):

If \[\text{ -3x-4y=7 }\] and \[\ -2x+4y=3\] are true, then \[\text{ -6x-8y=14 }\] and \[\ -6x+12y=9\] are also true (multiplied by 2 and 3 respectively)

OpenStudy (anonymous):

Therefore \[\ -6x=8y+14\] and \[\ -6x=-12y+9\] (re-arranging). So: \[\ -12y+9=8y+14\] You can find y using normal algebra now. Insert the found value for y into one of the first 2 equations, then find x.

OpenStudy (anonymous):

Or you can add the two equations together and get -5x=10.

OpenStudy (anonymous):

henpen's method is correct, but needlessly complicated and is not an example of the elimination method; it is an example of substitution via multiplication and the reflexive property/transitive property.